OpenStudy (youngnephew34):

What are the values of a and b? A) A=9/2, B=15/2 B) A=15/2, B=9/2 C) A= 16/3, B= 15/2 D) A=9/2, B= 13/2

OpenStudy (anonymous):

Is thee a pic i can refer too?

OpenStudy (youngnephew34):

I'm trying to get it

OpenStudy (anonymous):

Ill help you out when you have the picture

OpenStudy (anonymous):

Tag me when you get it?

OpenStudy (youngnephew34):

@CallMeEmmie

OpenStudy (youngnephew34):

I think it's A

OpenStudy (anonymous):

what grade are you in

OpenStudy (youngnephew34):

11

OpenStudy (anonymous):

I cant help than I am only in 6ths

jimthompson5910 (jim_thompson5910):

Hint: The two smaller triangles are similar triangles. So we can form this proportion a/6 = 6/8

OpenStudy (youngnephew34):

ok

jimthompson5910 (jim_thompson5910):

tell me what you get when you solve for 'a'

OpenStudy (youngnephew34):

I did it different

jimthompson5910 (jim_thompson5910):

how did you do it?

OpenStudy (youngnephew34):

I did 18/a = a/8

OpenStudy (youngnephew34):

and I'm still confuse

OpenStudy (youngnephew34):

and I also did like 18/b=b/8 and I got 72

jimthompson5910 (jim_thompson5910):

you can't add the 10 and 8 to get 18

OpenStudy (youngnephew34):

oh ok

jimthompson5910 (jim_thompson5910):

when you solve `a/6 = 6/8` for 'a', what did you get?

OpenStudy (youngnephew34):

I think you suppose to find the common denominator

OpenStudy (youngnephew34):

so 2in to 6 will be 3 and 2 goes into 8 4

jimthompson5910 (jim_thompson5910):

why not multiply both sides by 6 ?

jimthompson5910 (jim_thompson5910):

try that and see what you get

OpenStudy (youngnephew34):

8a=36

jimthompson5910 (jim_thompson5910):

now divide both sides of `8a = 36` by 8 to isolate 'a'

OpenStudy (youngnephew34):

36/8=4.5

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

a = 4.5 or a = 9/2

jimthompson5910 (jim_thompson5910):

now use that value of 'a' and the pythagorean theorem to find 'b'

OpenStudy (youngnephew34):

4.5x4.5=20.5

jimthompson5910 (jim_thompson5910):

that's the value of a^2, yes

OpenStudy (youngnephew34):

ok

jimthompson5910 (jim_thompson5910):

one of your steps will have a^2 + 6^2 = b^2 solve for b

OpenStudy (youngnephew34):

6x6=36

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (youngnephew34):

ok

jimthompson5910 (jim_thompson5910):

a^2 + 6^2 = b^2 turns into 20.25 + 36 = b^2

OpenStudy (youngnephew34):

s0 56.25

jimthompson5910 (jim_thompson5910):

that's the value of b^2 but we want the value of b

OpenStudy (youngnephew34):

so find the square root of 56.25

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (youngnephew34):

7.5

jimthompson5910 (jim_thompson5910):

correct, b = 7.5

jimthompson5910 (jim_thompson5910):

or b = 15/2

OpenStudy (youngnephew34):

oh ok

OpenStudy (anonymous):

lol ya'll are doing it wrong. im in 12th grade..this pellet is easy. ur gonna get it wrong

jimthompson5910 (jim_thompson5910):

since 15/2 = 7.5

OpenStudy (youngnephew34):

how

OpenStudy (youngnephew34):

and ok

OpenStudy (youngnephew34):

I'm still here

OpenStudy (youngnephew34):

so it's B

jimthompson5910 (jim_thompson5910):

Summary: a = 4.5 or a = 9/2 b = 7.5 or b = 15/2

OpenStudy (youngnephew34):

B should be my correct answer